# Exact treatment of the quantum Langevin equation under time-dependent system-bath coupling via a train of delta distributions

### v1

@@ IV.1.1 Introducing the spectral density
and thereby the dissipation kernel $\Sigma(t,t^{\prime})$ reads
@@
By introducing the spectral density $\sigma(\omega)$, each term in the two-time correlation function can be rewritten in terms of it. This approach allows us to examine a variety of environments, including those with an infinite number of quantum harmonic oscillators. The idea is as follows. So far, our $\sigma(\omega)$ given in (40) is the explicit form for a finite collection of quantum harmonic oscillators. To consider other types of environments (e.g., the case where $\mathcal{N}\to\infty$), we simply replace the spectral density in (40) with the one of our interest. A widely examined spectral density is the Lorentz-Drude (LD) spectral density:
$\sigma(\omega)$ $=\gamma\dfrac{\omega\Lambda^{2}}{\omega^{2}+\Lambda^{2}}\,,$ (45)
where $\gamma>0$ is the coupling constant and $\Lambda>0$ is the cutoff frequency that determines the bandwidth of the environment. Moreover, $\Lambda^{-1}$ is the characteristic time scale of the change in the environment. Given a cutoff $\Lambda$, the low-frequency regime $\omega\ll\Lambda$ of the LD spectral density mimics the Ohmic system as $\sigma(\omega)\approx\gamma\omega$, which reflects the Markovian dynamics. Employing the LD spectral density, we have
$\Gamma(t)$ $=\gamma\Lambda^{2}e^{-\Lambda|t|}\text{sgn}(t)\,.$ (46)
IV.1.2 Spectral density for discretized data
When considering continuous (typically constant) switching functions $\chi(t)$, one can compute correlation functions using the spectral density defined above. In this paper, however, the train of delta switchings requires us to adopt the discrete-time Fourier transform (DTFT) [22]—the discrete-time variant of the continuous-time Fourier transform—to define the spectral density.
To see this, consider the two-time correlation functions such as $\braket{Q(t)Q(t^{\prime})}$. As we saw in (31), these correlation functions are composed of the terms such as $\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}$, which is the noise correlation function evaluated at the delta-switched (discrete) times $t_{k}$ and $t_{k^{\prime}}$. Therefore, instead of employing the continuous-time Fourier transform to define the spectral density and the noise correlations, it is suitable to choose the DTFT for our delta-switched scenario. [^4] Simply inserting $t=t_{k}$ and $t^{\prime}=t_{k^{\prime}}$ in the noise correlator (44c) defined using the continuous-time Fourier transform leads to divergence at $k=k^{\prime}$.
@@ IV.1.2 Spectral density for discretized data
Figure 3: Comparison between the solutions to the QLE with a constant switching (dashed line) and our train of delta switchings (solid curve). The horizontal axis is taken to be the number of delta switchings $N$. Here, we choose $\gamma/\Omega=0.1,\Lambda/\Omega=2$. (a) The function $G_{f_{P}}$ at fixed time $T=1$ asymptotes to $G(t=1)$ as the number of delta switchings $N$ increases. (b) The correlation function evaluated with the train of delta switchings also asymptotes to the result for constant switching.
@@
$\Gamma_{k}$ $=\gamma\Lambda^{2}e^{-\Lambda T|k|/N}\text{sgn}(k)\,.$ (50)
@@
$\breve{\Gamma}(\overline{\omega})$ $=\sum_{k=-\infty}^{\infty}\Gamma_{k}e^{\mathsf{i}\overline{\omega}k}=\dfrac{-\mathsf{i}\gamma\Lambda^{2}\sin\overline{\omega}}{\cos\overline{\omega}-\cosh(\Lambda T/N)}\,,$ (51)
@@
$s(\overline{\omega})$ $=-\dfrac{1}{2}\dfrac{\gamma\Lambda^{2}\sin\overline{\omega}}{\cos\overline{\omega}-\cosh(\Lambda T/N)}\,.$ (52)
@@
IV.2 Comparison to constant switching
@@
Figure 3(a) shows $G_{f_{P}}(T)$ evaluated against the number of delta switchings $N$. Here, we choose $T=1$. As the number of switching times $N$ increases, our $G_{f_{P}}(T)$ asymptotes to the well-known continuous solution $G(t)$ at $t=1$. This indicates that the delta-switching method well-approximates the solution derived with the continuous switching. We also evaluate an element of covariance matrix $\braket{Q^{2}(t)}$ in Fig. 3(b) when the system is initially prepared in a coherent state. As $G(t)$ is already well-approximated by $G_{f_{P}}(t)$, the correlation function $\braket{Q^{2}(t)}$ is also approximated by delta switchings when $N$ is large enough.
@@ V Conclusion
We considered the QLE in the Caldeira-Leggett model with a time-dependent coupling and showed that a train of Dirac delta switchings is a powerful tool to deal with the equation. In particular, the delta switchings allow us (i) to solve the QLE exactly even if the dissipation kernel is not time-translation invariant, and (ii) to naturally incorporate the notion of discretized data (i.e., the samples of continuous functions) so that the DTFT can be employed for computing two-point correlation functions.
The solution to the QLE using the delta switchings is expressed as a sum of all possible influences of the memory effect in an environment. This can be intuitively understood by using diagrams in Figs. 1 and 2, which can also help us to construct the solution $Q(t)$ without actually solving the equation.
Our method can also be applied to the QLE in the relativistic settings. In this case, the system is modeled by a harmonic oscillator-type Unruh-DeWitt particle detector [23, 24] coupled to a quantum field in curved spacetime. However, it is crucial to introduce a smearing function (i.e., a detector’s size) to avoid the UV divergences due to the nature of delta-switchings. This is currently under investigation by the authors.
@@ Acknowledgements.
K.G.Y. is supported by Grant-in-Aid for JSPS Fellows Grant Number JP25KJ0048. TT gratefully acknowledges the support of K2-Spring for providing the resources necessary for this research.
Appendix A Markovianity and the covariance matrix
A.1 Transformation of covariance matrices

### v3

@@ IV.3.1 Introducing the spectral density
and thereby the memory kernel $\Sigma(t,t^{\prime})$ reads
@@
By introducing the spectral density $\sigma(\omega)$, each term in the two-time correlation function $\braket{R_{i}(t)R_{j}(t^{\prime})}$ can be rewritten in terms of it. This approach allows us to examine a variety of environments, including those with an infinite number of quantum harmonic oscillators. The idea is as follows. So far, our $\sigma(\omega)$ given in (70) is the explicit form for a finite collection of quantum harmonic oscillators. To consider other types of environments (e.g., the case where $\mathcal{N}\to\infty$), we simply replace the spectral density in (70) with the one of our interest. A widely examined spectral density is the Lorentz–Drude (LD) spectral density:
$\sigma(\omega)$ $=\kappa\dfrac{\omega\Lambda^{2}}{\omega^{2}+\Lambda^{2}}\,$, (75)
where $\kappa>0$ is the coupling constant and $\Lambda>0$ is the cutoff frequency that determines the bandwidth of the environment. Moreover, $\Lambda^{-1}$ is the characteristic time scale of the change in the environment. Given a cutoff $\Lambda$, the low-frequency regime $\omega\ll\Lambda$ of the LD spectral density mimics the Ohmic system as $\sigma(\omega)\approx\kappa\omega$, which reflects the Markovian dynamics. Employing the LD spectral density, we have
$\Gamma(t)$ $=\kappa\Lambda^{2}e^{-\Lambda|t|}\text{sgn}(t)\,.$ (76)
IV.3.2 Spectral density for discretized data
When considering continuous (typically constant) switching functions $\chi(t)$, one can compute correlation functions $\braket{R_{i}(t)R_{j}(t^{\prime})}$ using the spectral density defined above. In this paper, however, the train of delta switchings requires us to adopt the discrete-time Fourier transform (DTFT) [24]—the discrete-time variant of the continuous-time Fourier transform—to define the spectral density.
To see this, consider the two-time correlation functions such as $\braket{Q(t)Q(t^{\prime})}$. As we saw in (61), these correlation functions are composed of the terms such as $\braket{\zeta(t_{k})\zeta(t_{k^{\prime}})}$, which is the noise correlation function evaluated at the delta-switched (discrete) times $t_{k}$ and $t_{k^{\prime}}$. This suggests that one deals with sequences instead of continuous functions to compute the correlation functions. Therefore, instead of employing the continuous-time Fourier transform to define the spectral density and the noise correlations, it is suitable to choose the DTFT for our delta-switched scenario. [^6] Simply inserting $t=t_{k}$ and $t^{\prime}=t_{k^{\prime}}$ in the noise correlator (74c) defined using the continuous-time Fourier transform leads to divergence at $k=k^{\prime}$.
@@ IV.3.2 Spectral density for discretized data
Figure 5: Comparison between the solutions to the QLE with a constant switching (dashed line) and our train of delta switchings (solid curve). The horizontal axis is taken to be the number of delta switchings $N$. Here, we choose $\kappa/\Omega=0.1,\Lambda/\Omega=2$. (a) The function $G_{f_{P}}$ at fixed time $T=1$ (in units of $\Omega$) asymptotes to $G(t=1)$ as the number of delta switchings NN increases. (b) The correlation function evaluated with the train of delta switchings also asymptotes to the result for constant switching.
@@
$\Gamma_{k}$ $=\kappa\Lambda^{2}e^{-\Lambda T|k|/N}\text{sgn}(k)\,.$ (80)
@@
$\breve{\Gamma}(\overline{\omega})$ $=\sum_{k=-\infty}^{\infty}\Gamma_{k}e^{\mathsf{i}\overline{\omega}k}=\dfrac{-\mathsf{i}\kappa\Lambda^{2}\sin\overline{\omega}}{\cos\overline{\omega}-\cosh(\Lambda T/N)}\,$, (81)
@@
$s(\overline{\omega})$ $=-\dfrac{1}{2}\dfrac{\kappa\Lambda^{2}\sin\overline{\omega}}{\cos\overline{\omega}-\cosh(\Lambda T/N)}\,.$ (82)
@@
IV.4 Comparison to constant switching
@@
Figure 5(a) shows $G_{f_{P}}(T)$ evaluated against the total number of delta switchings $N$. Here, we choose $T=1$. As the number of switching times $N$ increases, our $G_{f_{P}}(T)$ asymptotes to the well-known continuous solution $G(t)$ at $t=1$. This indicates that the delta-switching method well-approximates the solution derived with the continuous switching. We also evaluate an element of covariance matrix $\braket{Q^{2}(t)}$ in Fig. 5(b) when the system is initially prepared in a coherent state. As $G(t)$ is already well-approximated by $G_{f_{P}}(t)$, the correlation function $\braket{Q^{2}(t)}$ is also approximated by delta switchings when $N$ is large enough.
@@ V Conclusion
By employing the train of delta-switchings method, we analytically solved a generic inhomogeneous integro-differential equation. The advantage of this method is that it is applicable to those with a nonstationary memory kernel, and that it allows us to visualize the memory effect using diagrams in Figs. 1 and 2, which can also help us to construct the solution without actually solving the equation. We applied our solution (29) to the Jaynes–Cummings model and the Caldeira–Leggett model. We found that our method well-approximates the well-known exact solutions in the continuum limit, and that the memory propagation between adjacent times corresponds to Markovian behavior.
Our method can also be applied to quantum systems in the relativistic settings (see the recent development in [2]). In this case, the system is modeled by (a qubit-type or a harmonic oscillator-type) Unruh-DeWitt particle detector [29, 14] coupled to a quantum field in curved spacetime. However, it is crucial to introduce a smearing function (i.e., a detector’s size) to avoid the UV divergences due to the nature of delta-switchings. This is currently under investigation by the authors.
@@ Acknowledgements.
We thank the anonymous referees for constructive comments and suggestions. K. G.-Y. was partially supported by Grant-in-Aid for Research Activity Start-up (Grant No. JP24K22862) and by Grant-in-Aid for JSPS Fellows (Grant No. JP25KJ0048). T. T. was supported by R7 (2025) Young Researchers Support Project, Faculty of Science, KYUSHU UNIVERSITY. T. T. gratefully acknowledges the support of K2SPRING for providing the resources necessary for this research.
Appendix A Review of Markovian dynamical maps
In this section, we introduce the concept of Markovian dynamical maps. The quantum Markovianity of a quantum dynamical map is often characterized by the completely positive trace-preserving (CPTP) property, along with the phenomenologically motivated semigroup property, which ensures an irreversible dynamical process. These properties are equivalent to the dynamical map satisfying the well-known GKSL master equation [5], [^8] The CPTP and semigroup properties are equivalent to the generalized GKSL master equation, where the Hamiltonian, coupling, and Kraus operators can depend on time. forming the foundation of Markovian quantum dynamics.
In this work, we define Markovian dynamics as those that can be expressed as a CPTP semigroup map. However, we emphasize that this is not the only definition of Markovianity, as various alternative characterizations exist [20].
Let us begin by defining Markovianity [27]. Consider a dynamical map that maps a system’s arbitrary quantum state from $t=0$ to time $t$, $\mathcal{E}_{t,0}:\rho_{\text{s}}(0)\mapsto\rho_{\text{s}}(t)$. Dynamical maps can be utilized even when a system is initially entangled with an environment. A completely-positive and trace-preserving (CPTP) map is a special class of dynamical maps, where the system’s input state $\rho_{\text{s}}(0)$ is assumed to be uncorrelated with the environment. For example, a map induced by $\rho_{\text{s}}(t)=\operatorname{Tr}_{\text{E}}[U_{t}\rho_{\text{tot}}(0)U_{t}^{\dagger}]$ is a CPTP if the system is initially uncorrelated with the environment.
A Markovian dynamics can be characterized in terms of the divisibility of a CPTP map. Suppose a dynamical map $\mathcal{E}_{t,0}$ is a CPTP map and it can be decomposed into two maps:
In general, however, it is not necessarily true that both $\mathcal{E}_{t,s}$ and $\mathcal{E}_{s,0}$ are CPTP maps. For example, $\mathcal{E}_{s,0}$ can be a CPTP map, but $\mathcal{E}_{t,s}$ is generally not, as the first CPTP map $\mathcal{E}_{s,0}$ entangles the system and the environment. A CPTP map $\mathcal{E}_{t,0}$ is called a Markovian dynamical map if both dynamical maps $\mathcal{E}_{t,s}$ and $\mathcal{E}_{s,0}$ are CPTP maps. This is also known as the one-parameter semigroup property, and it can be shown that it is equivalent to the generalized GKSL equation [27].
Appendix B Proof of Markovianity in the damped Jaynes–Cummings Model
As discussed in Sec. III.2, the semigroup property of the quantum channel reduces to a property of the time-evolution function of the probability amplitude, given by Eq. (37), as follows:
$\mathcal{T}(t)$ $=\mathcal{T}(t,t^{\prime})\mathcal{T}(t^{\prime})\,$, (87)
for any $t^{\prime}\in[0,t)$. Below, we prove that our solution (37) satisfies this Markovianity condition if the memory propagation in Fig. 2 is restricted to nearest-neighbor propagation $t_{l}\xrightarrow{\Sigma}t_{l+1}$, where $t_{l}\equiv lT/N$. In practice, we impose the Markov approximation $\Lambda T/N\gg 1$ to ensure that the nearest-neighbor propagation is dominant. To this end, consider $\mathcal{T}(t)$ in (37) only composed of the neighboring propagation $\Sigma(t_{l+1},t_{l})$, which reads
$\mathcal{T}(t,0)$ $=1+\sum_{i_{1}=1}^{N-1}X_{i_{1}}+\sum_{i_{1}=1}^{N-1}\sum_{i_{2}=1}^{i_{1}-1}X_{i_{1}}X_{i_{2}}+\cdots$
$=\mathsf{T}\exp\left(\sum_{j=1}^{N-1}X_{j}\right),$ (88)
where
$X_{j}\equiv-\dfrac{T^{2}}{N^{2}}\Sigma(t_{j+1},t_{j})\,$, (89)
and $\mathsf{T}$ denotes the time-ordering symbol. Note that we take $t=t_{N}$. Based on the above expression, the time-evolution function $\mathcal{T}(t_{b},t_{a})$ from time $t_{a}$ to $t_{b}$, with integers $a<b(<N)$, can be written as
$\mathcal{T}(t_{b},t_{a})=\mathsf{T}\exp\left(\sum_{j=a}^{b-1}X_{j}\right)\,.$
We now proceed to verify the semigroup property (87). Since $t>t^{\prime}>0$, without loss of generality, let us denote $t^{\prime}=t_{a}$ for any integer $a\in(1,N)$, and $t=t_{N}$. Then we find that
$\mathcal{T}(t,t_{a})\mathcal{T}(t_{a})$ $=\mathsf{T}\exp\left(\sum_{j=a}^{N-1}X_{j}\right)\mathsf{T}\exp\left(\sum_{j=1}^{a-1}X_{j}\right)$
$=\mathsf{T}\exp\left(\sum_{j=1}^{N-1}X_{j}\right)=\mathcal{T}(t)\,$, (90)
which completes our proof.
Appendix C Measures of non-Markovianity
In this appendix, we show that when $\kappa T/N$ is finite, the $g$-function (defined below) used in the RHP measure equals the decay rate plus a correction term. This correction term vanishes in the limit $\kappa T/N\rightarrow 0$, rendering a continuous switching case.
First, let us state the RHP measure introduced in Ref. [26]. Let $\mathcal{E}_{t,0}:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H})$ be a Markovian dynamical map. Let us consider an infinitesimal evolution beyond the final time $t$ with $\mathcal{E}_{t+\epsilon,t}$. Then a composite map is given by
$\mathcal{E}_{t+\epsilon,0}=\mathcal{E}_{t+\epsilon,t}\circ\mathcal{E}_{t,0}.$ (91)
If the infinitesimal increment $\mathcal{E}_{t+\epsilon,t}$ is a CPTP map, then $\mathcal{E}_{t+\epsilon,0}$ is also a CPTP map. This small incremental evolution can be done many times to construct any time evolution beyond $t$, which ensures the semi-group property:
$\mathcal{E}_{t_{1},t_{2}}=\mathcal{E}_{t_{1},s}\circ\mathcal{E}_{s,t_{2}},$ (92)
where $t_{1}\geq s\geq t_{2}\geq 0$.
The CPTP property of the dynamical map $\mathcal{E}_{t+\epsilon,t}$ is ensured by showing that the Choi matrix of the map $\mathcal{E}_{t+\epsilon,t}$ is positive definite, which is true if the trace-norm of the Choi matrix is equal to 1:
$\|(\mathcal{E}_{t+\epsilon,t}\otimes\mathds{1}_{\text{A}})\left[\Psi^{+}_{\text{SA}}\right]\|=1$ $\text{ iff }\mathcal{E}_{t+\epsilon,t}\text{ is CP,}$ (93a)
$\|(\mathcal{E}_{t+\epsilon,t}\otimes\mathds{1}_{\text{A}})\left[\Psi^{+}_{\text{SA}}\right]\|>1$ $\text{ iff }\mathcal{E}_{t+\epsilon,t}\text{ is non-CP,}$ (93b)
[…]
Using the above relation authors in Ref. [26] constructed the $g$-function:
$g(t)\coloneqq\lim_{\epsilon\rightarrow 0^{+}}\dfrac{\|(\mathcal{E}_{t+\epsilon,t}\otimes\mathds{1}_{\text{A}})\left[\Psi^{+}_{\text{SA}}\right]\|-1}{\epsilon},$ (94)
which gives zero if and only if the map $\mathcal{E}_{t+\epsilon,t}$ is CP, and becomes positive if and only if $\mathcal{E}_{t+\epsilon,t}$ is non-CP. Integrating this function over all time scales, we obtain the RHP measure:
$\mathcal{I}\coloneqq\int_{0}^{\infty}\text{d}t\,g(t),$ (95)
which can be used as a measure of Markovianity: $\mathcal{I}=0$ indicating the Markovianity, and $\mathcal{I}\neq 0$ indicating the non-Markovianity. In the damped Jaynes-Cumming model the $g$-function is related to the decay rate $\gamma(t)$ of the system in the following way
$g(t)=\left\{\begin{array}[]{cc}0&\text{ if }\gamma(t)\geq 0\\ -\gamma(t)&\text{ if }\gamma(t)<0\end{array}\right.\,$. (98)
We will show below that this relation only holds in our Dirac-delta switching method when $\kappa T/N$ tends to zero.
To find the relation between the $g$-function and the decay rate $\gamma(t)$ in our delta-switching scenario, we first need to obtain the time-convolutionless master equation for our setup. Following the method in Ref. [5], we take the time derivative of the system density matrix at an arbitrary time $t$, given in Eq. (39):
[…]
$=\begin{bmatrix}2\dot{\alpha}_{1}(t)\alpha_{1}^{*}(t)&\alpha_{0}^{*}\dot{\alpha}_{1}(t)\\ \alpha_{0}\dot{\alpha}_{1}^{*}(t)&-2\dot{\alpha}_{1}(t)\alpha_{1}^{*}(t)\end{bmatrix}+\begin{bmatrix}\delta|\dot{\alpha}_{1}(t)|^{2}&0\\ 0&-\delta|\dot{\alpha}_{1}(t)|^{2}\end{bmatrix}\,$, (99)
where $\delta\equiv T/N$ is the time interval between Dirac-delta switching. We have also defined a quantity $\dot{\alpha}_{1}(t)\equiv(\alpha_{1}(t+\delta)-\alpha_{1}(t))/\delta$. Rewriting this equation in terms of the Pauli matrices $\sigma_{\pm}=(\sigma_{x}\pm\mathsf{i}\sigma_{y})$, we find
$\dfrac{\rho_{\text{s}}(t+\delta)-\rho_{\text{s}}(t)}{\delta}=$ $\gamma(t)\Big[\sigma_{+}\rho_{\text{s}}(t)\sigma_{-}-\dfrac{1}{2}\{\sigma_{-}\sigma_{+},\rho_{\text{s}}(t)\}\Big]$
$+\delta h(t)\operatorname{Tr}[\rho_{\text{s}}(t)][\sigma_{-},\sigma_{+}]$
$\equiv$ $\mathcal{L}_{t}[\rho_{\text{s}}(t)]\,$, (100)
where we have defined the two functions
$\gamma(t)\equiv-2\text{Re}\left(\frac{\dot{\alpha_{1}}(t)}{\alpha_{1}(t)}\right),\quad h(t)\equiv\left|\dot{\alpha_{1}}(t)\right|^{2},$ (101)
where $\gamma(t)$ corresponds to the decay rate of our system. Notice that if we take $\delta|\dot{\alpha}_{1}(t)|^{2}\rightarrow 0$ in the above equation, we obtain the usual time-convolutionless master equation for the damped Jaynes-Cummings model Ref. [5].
Rewriting the left-hand side of the above equation with $\rho_{\text{s}}(t+\delta)=\mathcal{E}_{t+\delta,t}[\rho_{\text{s}}(t)]$, we find
$\mathcal{E}_{t+\delta,t}[\rho_{\text{s}}(t)]=(\mathds{1}_{\text{S}}+\mathcal{L}_{t}\delta)[\rho_{\text{s}}(t)]\,.$ (102)
Finally, the Choi matrix of the above quantum channel gives a $4\times 4$ matrix of the following form:
$\mathds{1}_{\text{SA}}+\delta(\mathcal{L}_{t}\otimes\mathds{1}_{\text{A}}))[\Psi_{\text{SA}}^{+}]$
$=\dfrac{1}{2}\begin{bmatrix}0&\delta^{2}h(t)&0&0\\ 0&1-\delta\gamma(t)+\delta^{2}h(t)&1-\delta\frac{\gamma(t)}{2}&0\\ 0&1-\delta\frac{\gamma(t)}{2}&1&-\delta^{2}h(t)\\ 0&0&0&\delta\gamma(t)-\delta^{2}h(t)\end{bmatrix}\,$, (103)
with the following four eigenvalues:
$\lambda_{0}=0,~\lambda_{1}=\frac{\delta}{2}(\gamma(t)-\delta h(t)),$ (104)
$\lambda_{\pm}=\frac{2-\delta(\gamma(t)-\delta h(t))}{4}$
$~\pm\frac{\sqrt{4-4\delta\gamma(t)+2\delta^{2}\gamma(t)^{2}-2\delta^{3}h(t)\gamma(t)+\delta^{4}h(t)^{2}}}{4}.$ (105)
To calculate the trace-norm, we take the absolute value sum of all four eigenvalues. However, since our interest lies in the small time interval $T/N$, we can use the perturbation theory. By inspection, notice that eigenvalues depend on the combination $\delta^{2}h(t)=\mathcal{O}(\kappa T/N)$, where $\kappa$ is the coupling strength between the system and the environment. Therefore, we can Taylor expand the trace-norm with respect to the dimensionless parameter $\kappa T/N$:
$\|(\mathds{1}_{\text{SA}}+\delta(\mathcal{L}_{t}\otimes\mathds{1}_{\text{A}}))[\Psi_{\text{SA}}^{+}]\|$
$\quad=\left|1-\dfrac{1}{2}\delta\gamma(t)\right|+\dfrac{\delta}{2}|\gamma(t)|+\mathcal{O}((\kappa\delta)^{2})$
$\quad=1-\dfrac{1}{2}\delta\gamma(t)+\dfrac{\delta}{2}|\gamma(t)|+\mathcal{O}((\kappa\delta)^{2}),$ (106)
where we observe that there is a correction term of order $(\kappa\delta)^{2}=(\kappa T/N)^{2}$. Calculating the $g$-function, we find
$g(t)=\begin{cases}0+\mathcal{O}(\kappa^{2}\delta)&\text{if}\quad\gamma(t)\geq 0\\ -\gamma(t)+\mathcal{O}(\kappa^{2}\delta)&\text{if}\quad\gamma(t)<0\end{cases}.$ (107)
Since this is a small correction term, we take the decay rate $\gamma(t)$ to calculate the RHP measure in Sec. III. However, we note that in the strong coupling limit or the long-time interval case, the correction term $\kappa T/N$ cannot be ignored, and the decay rate is no longer the measure of Markovianity. It is interesting to study such cases, but since this is outside the scope of this paper, we leave it for future work.
Appendix D Markovianity and the covariance matrix
D.1 Transformation of covariance matrices